Insoluble Li2Co3 (s) precipitates when solutions of Na2Co3 (aq) and LiCl (aq) are mixed.

To measure the enthalpy change, 250. mL of 0.80 M Na2Co3 (aq) and 250. mL of 1.6 M LiCl (aq) are mixed in a coffee-cup calorimeter. The temperature of the mixture rises by 1.6 °C. Calculate the enthalpy change for the precipitation of Li2Co3 (s) in kJ/mol. (Assume the density of the solution is 1.0 g/mL, and its specific heat capacity is 4.2 J/g · K.)
Answer #1

Solution-

(4.2 J/g·K) x (250 mL + 250 mL) x (1.0 g/mL) x (1.6 K) = 3360 J

(0.250 L) x (1.6 mol/L LiCl) x (1 mol Li2CO3 / 2 mol LiCl) = 0.2 mol of Li2CO3

(3360 J)/( 0.2 mol of Li2CO3) = 16800 J/mol = 16.8 kJ/mol of Li2CO3

Here the enthalpy is negative by convention because the reaction was exothermic

Hence -16.8 kJ/mol Li2CO3